# A Survey of the Self Supervised Learning Mechanisms for Vision Transformers

### v1

@@ Abstract
Deep supervised learning models require high volume of labeled data to attain sufficiently good results. Although, the practice of gathering and annotating such big data is costly and laborious. Recently, the application of self supervised learning (SSL) in vision tasks has gained significant attention. The intuition behind SSL is to exploit the synchronous relationships within the data as a form of self-supervision, which can be versatile. In the current big data era, most of the data is unlabeled, and the success of SSL thus relies in finding ways to improve this vast amount of unlabeled data available. Thus its better for deep learning algorithms to reduce reliance on human supervision and instead focus on self-supervision based on the inherent relationships within the data. With the advent of ViTs, which have achieved remarkable results in computer vision, it is crucial to explore and understand the various SSL mechanisms employed for training these models specifically in scenarios where there is less label data available. In this survey we thus develop a comprehensive taxonomy of systematically classifying the SSL techniques based upon their representations and pre-training tasks being applied. Additionally, we discuss the motivations behind SSL, review popular pre-training tasks, and highlight the challenges and advancements in this field. Furthermore, we present a comparative analysis of different SSL methods, evaluate their strengths and limitations, and identify potential avenues for future research.
@@ 1 Introduction
To tackle above-mentioned challenges in supervised learning, several learning strategies are presented, such as active learning [6], semi-supervised learning [7], and SSL [8]. Recently, Transformers [9] have emerged as an effective Neural Network (NN) architecture for leveraging deeper insights in the input Data, specifically, the ViTs [10] are considered to be the alternative of CNNs [11]. ViTs have been successfully employed in tasks related to computer vision, including object detection, recognition and semantic segmentation. Notably, ViTs have demonstrated superior performance in image classification, when trained with large volume datasets such as JFT-300M [12].
In essence, by designing pretext tasks, a pretraining phase of network that encourage the model to learn significant patterns in the data itself, SSL methods enable ViTs to understand underlying structures of data without any supervision. The motivation for applying SSL in ViTs is two-fold. Firstly, SSL allows ViTs to employ the vast amount of unlabeled data that is readily available, allowing them to develop more robust and generalizable representations. SSL is beneficial in situations where labeled data is scarce or costly to acquire. Secondly, SSL provides a pathway for pretraining ViTs on larger volume of datasets, followed by process of fine-tuning on specific tasks, leading to improved performance and faster convergence [5].
@@ 1.2 Survey Structure
The paper begins with an introduction to SSL and a detailed discussion on importance of SSL in ViTs. The reported SSL methods are segmented into 5 classes according to their unique features. The paper discuss the recent advancements and applications of SSL across different computer vision tasks. It also presents a taxonomy for SSL, which classifies them based on how feature learning is applied within their architecture. According to this taxonomy, SSLs are divided into five groups, each representing unique way of making use of input features. Frequently used abbreviations are listed in Table 1. The structure of paper is depicted in Fig 1. Section 1 discuss a systematic understanding of the SSL architecture, highlighting its need in ViTs and outlined the advent of SSL architectures. Proceeding to section 2 which covers the advancements in SSL variants, while section 3 provide a taxonomy of the recent SSL architectures, respectively. Section 4 focuses on the regularization techniques used in SSL, particularly in the area of computer vision, section 5 presents different metrics and benchmarks used to examine the effectiveness of SSL while section 6 discusses comparative analysis of state-of-the-art SSL mechanisms employing vision transformers, section 7 mentions some open challenges in domain of SSL, section 8 discusses different applications and future directions of SSL for vision transformers current challenges and potential developments. Finally section 9 concludes this survey paper.
@@
CLIP Contrastive Language-Image Pretraining
@@ 2 Evolution of SSL
In computer vision SSL methods are generally segmented into contrastive, generative, and predictive approaches. Contrastive methods, such as MoCo and SimCLR, aim to learn patterns by contrasting positive and negative samples. Generative methods, such as GANs and VAEs, focus on learning representations by generating samples from the learned representation space. Predictive methods, such as BERT and T5, learn representations by predicting some parts of the input data.
one SSL category is instance discrimination, in which the motto of learning is basically focused on learning patterns by differentiating each sample image from others, and this method is rather challenging for large volume of datasets. Examples are SimCLR, SwAV, BYOL, MoCO [18, 19, 20, 21].
@@ 2.1 Evolution of ViTs
In computer vision, the traditional approach in image processing tasks is the utilization convolutional neural networks (CNNs). However, the CNNs have limitations in capturing dependencies across long spatial distances and global context, while both are crucial for some vision tasks. ViTs address these limitations via self-attention mechanisms to learn and understand the global context and also dependencies over long spatial distances in the input data .
The initial vision transformer architecture ViT, was presented in the paper ”An Image Is Worth 16x16 Words” [10] Since then, various variants of ViTs are proposed, including DeiT, Swin Transformer, and TNT. These architectures have obtained state-of-the-art performance in different computer vision based tasks, including image classification, object segmentation and object detection.
@@ 2.2 Importance of SSL in ViTs
SSL is especially crucial in the context of ViTs, as it empowers models to learn representations from large amount of unlabeled data. ViTs need vast amounts of data to learn effective representations, and SSL provides a pathway for pretraining on large-scale datasets. By pretraining on large-scale datasets, ViTs can learn more robust.
@@ 2.3 Evolution of SSL in Computer Vision
SSL technique is used for representation learning where a model is designed to learns the underlying relationships between its inputs. The framework of Energy-Based Models (EBMs) can effectively describe this objective, where the aim of assigning high energy to inputs that are incompatible and low energy to those that are compatible. The foundation for many self-supervised learning techniques applied to images was laid by Geoffrey Hinton and others in the 1980s through the development of autoencoders and the proposal of greedy layer-wise pretraining [22], where layers of a deep neural networks are trained sequentially. However, autoencoders were basically used for learning features and reducing dimensionality. They were impactful at that time as they made it possible to train the first “deep” networks. An analogous approach was Restricted Boltzmann Machines (RBMs), that allows layer-wise training and then combine those layers to build deep belief nets [23]. Although these techniques were eventually replaced by simpler initialization strategies and extended training procedures, they played a vital part in the development of early deep learning architectures.
@@
In 2000s, types of unsupervised learning, including sparse coding and K-means, were employed for self-supervised learning (SSL). These techniques aimed to learn image features without labeled data, laying the groundwork for more complex self-supervised methods. The category of Spatial Context prediction gained attention in SSL for the training of CNNs. In this regard, Pathak et al. (2016) proposed inpainting using context prediction as a pretext task for SSL in images, where the model predicts missing parts of an image. Zhang et al. (2016) explored a straightforward method of the colorization task for self-supervised representation learning. They trained a model to colorize grayscale images, which significantly enhances its capability to understand textures, patterns, and object identities. Noroozi and Favaro (2016) [24] introduced a method where the model learns to solve jigsaw puzzles, using the relative positions of image patches as supervision. Gidaris et al. (2018) introduced RotaNet [25], which employs a pretext task where the model predicts the rotation angle applied to an image, enhancing feature learning for various orientations. Meanwhile, clustering approaches also progressed in SSL. Caron et al. (2018) proposed DeepCluster [26], a method that iteratively clusters image features and trains a neural network to predict the cluster assignment.
@@
Contrastive learning became significant approach in self-supervised learning during the era of convolutional neural networks (CNNs) (He et al., 2020; Chen et al., 2020). The roots of contrastive learning can be found in the early work on metric learning and the development of Siamese networks in the 1990s and early 2000s. Siamese Network architecture [27], introduced by Bromley et al. in 1993, involved training twin networks to lessen the relationship among embeddings of matching pairs and expanding the gap between embeddings of constrasting pairs. The formalization of contrastive loss by Hadsell, Chopra, and LeCun (2006) was a significant step forward [28].
@@
In the late 2010s, the concept of employing contrastive learning for self-supervised learning gained attraction. This period saw the development of several influential techniques that leveraged contrastive objectives to learn meaningful patterns from unlabeled data. Dosovitskiy et al. (2014) used instance discrimination as a pretext task. Each image was handled as a different class, and the network was trained to differentiate between differnt variations of the same image. Wu et al. (2018) refined the instance discrimination approach using a memory repository to save negative examples for the contrastive loss. Oord et al. (2018) introduced method known as Contrastive Predictive Coding in 2018, which learns patterns by predicting future observations in latent space using contrastive learning. Chen et al. (2020) introduced a framework known as SimCLR for contrastive learning of visual features that trains by contrasting similar and non-similar pairs of data. This significantly advanced the performance of self-supervised methods in computer vision. Over time, the idea of hybrid learning gained strength, and in 2020, Caron et al. (2020) introduced SwAV [19], which combines contrastive learning with clustering by using swapped assignments between different augmentations of an image.
@@
The idea of self-distillation give a new dimension to SSL. Grill et al. (2020) boosted the idea of SSL by introducing BYOL, a self-supervised method that avoids the need for negative samples (simplifying contrastive learning) by using two networks to predict each other’s outputs. He et al. (2020) introduced MoCo, which extends the idea of SimCLR by using a dynamic dictionary and a momentum encoder for contrastive learning. Facebook AI Research (Caron et al., 2021) presented DINO [29], which uses self-distillation with no labels, leveraging VITs for SSL.
@@
In 2020, the idea of transformers in vision was introduced. Vision transformers are high-capacity models and need a huge data for model tuning and good generalization. For improving the learning ability of the transformers, Dosovitskiy et al. (2020) adapted SSL techniques like masked image modeling and contrastive learning for ViT. Inspired by BERT (2018) [30], Dosovitskiy et al. masked certain segment tokens and replaced them with trained mask tokens. Later the model was further trained for predicting pixel values directly. However, they found unsupervised pre-training technique notably less useful than pre-training with supervised methods. Before ViTs, the same idea of inpainting was employed in CNNs, where portions of an image were masked out and the model was taught to reconstruct them, this approach to pre-training is termed as MIM (Masked Image Modeling). In ViTs, the MIM was framed as a regression task. This involved initially utilizing an autoencoder, which encode image segments to separate tokens, subsequently, the transformer is pre-trained to estimate the distinct token values for masked tokens. At that time, BEiT demonstrated substantial enhanced results in image classification and semantic segmentation compared to earlier supervised and self-supervised methods. However, its training process was intricate as it required a robust autoencoder to transform image patches into discrete tokens.
@@
Recently, the idea of multi-modality gained attention with the introduction of CLIP by OpenAI in 2021 [35]. CLIP learns image and text embeddings simultaneously by matching corresponding image-caption pairs. Similarly, ALIGN [36] uses self-supervised learning approaches to develop a shared feature space for images and their corresponding captions. The concept of MIM pre-training was streamlined by two concurrent works: MAE by He et al. (2022) and SimMIM by Xie et al. (2022). These approaches propose simplified algorithms that immediately rebuild masked parts of image patches instead of using separate image pieces obtained from an encoder as done in BEiT [37]. This idea was inspired by masked language modeling in NLP. These straightforward pre-training strategies performs better as compared to BEiT on classifying images, object detection and semantic segmentation (downstream tasks). MIM has recently dominated this domain by obtaining state-of-the-art performance on ViTs. The essence of this technique is to enhance the network’s capability to capture visual context at patch level using a denoising auto-encoding mechanism. Invariant Joint Embedding Predictive Architecture (IJEPA) [38] has emerged as a promising alternative, addressing some of the limitations inherent in other self-supervised learning methods. Unlike traditional self-supervised approaches that depend heavily on hand-crafted data augmentations, IJEPA focuses on invariant feature learning and joint embeddings. IJEPA relies on predicting the patterns of destination blocks within an image from a single context block, using a strategic masking method to guide the learning process.
@@
RotNet Spatial awareness, Object recognition Complex texture understanding
@@ 2.4 Comparison of SSL with Transfer Learning
In computer vision, transfer learning (TL) is employed in the development of CNN models to solve a new challenging problem by utilizing the pre-trained model’s weight sharing mechanism especially in the presence of data scarcity [39]. Initially, transfer learning was the dominant approach, but recently, SSL has demonstrated promising results across various applications.
@@ 3.0.1 Contrastive Methods
Contrastive Learning (CL) stands as a prominent technique in SSL (He et al., n.d.; Chen et al., 2020). Its aim is to capture invariant semantics through pairs of random views, termed Contrastive multi-view coding. CL ensures that global projections of representations are similar for positive samples and dissimilar for negative samples. This section delves into literature reviews on Contrastive Learning applied in transformer frameworks.
@@
Wang et al. (2022) introduced the method SRCL that is semantically-relevant contrastive learning for histopathology image analysis. Unlike traditional contrastive learning, SRCL evaluates similarity among instances to identify additional positive pairs by arranging various positive examples that share related visual concepts. This approach enriches the diversity of positive pairs and yields more informative patterns. They employee a hybrid model named CTransPath, in which they combines a CNN with a multi-scale Swin Transformer architecture, their methodology acts as a unified feature extractor, addressing both local and global features during pretraining, seeking to acquire global feature interpretations tailored for activities within the domain of histopathology images. The efficacy of SRCL-pretrained CTransPath was evaluated across five distinct downstream tasks spanning nine publicly available datasets, such as retrieving patches, classifying patches, performing weakly-supervised classification of whole-slide images, detecting mitosis, and segmenting glands in colorectal adenocarcinoma. The findings demonstrated that the visual representations derived from SRCL not just achieved state-of-the-art performance throughout all datasets but it also exhibited increased robustness and transferability compared to both supervised and self-supervised ImageNet pretraining techniques.
@@
To address instability issues during training, Chen et al. (2021) utilized a SSL framework in MOCO, representing an advancement in Contrastive Learning (CL) methodologies. MOCO utilizes dictionaries to gauge similarity between encoded keys and queries. Empirical assessments suggested that self-supervised Transformers within the MOCO framework can demonstrate good performance with relatively fewer inductive biases compared to ImageNet supervised ViT. This approach aims for a better balance between simplicity, accuracy, and scalability, showcasing its superiority over supervised pre-training methods in multiple detection and segmentation tasks. The study proposes replacing the patch projection layer in ViT with fixed random patch projections and suggests removing positional embeddings in SSL setups to address instability.
@@
Mo et al. (2023) introduced MCVT, which focuses on projecting class tokens to embedding space during the initial or final phases of the ViTs using multi-layer perceptrons. In this methodology, InfoNCE loss is applied to low-level features while ProtoNCE loss is introduced in high-level features, enhancing performance in downstream classification tasks. Extensive experimentation transferring MCVT pre-trained backbones to various downstream tasks confirmed its efficacy across multiple vision-related tasks.
@@
Radford et al. (2021) proposed a Contrastive Language-Image Pretraining (CLIP) framework, which incorporates transformers, learns image representations through natural language supervision. In pre-training, CLIP concurrently trains image and text encoders to accurately predict corresponding pairs within a batch of (image, text) training instances. CLIP discerns among the N * N possible (image, text) pairings within a batch. It achieves this by learning a multi-modal embedding space, optimizing the image and text encoders jointly to increase the cosine similarity of embeddings for the N genuine pairs while decreasing the resemblance of embeddings for the ( N 2̂ - N ) incorrect pairings. Symmetric cross-entropy loss is used in this optimization strategy to adjust these similarity scores. After pre-training, the model uses natural language to reference acquired visual concepts, facilitating zero-shot transfer to downstream tasks, demonstrating substantial transferability to diverse computer vision tasks without the need for task-specific training data. Evaluation encompasses benchmarking across more than 30 diverse computer vision datasets. Remarkably, the model demonstrates substantial transferability to most tasks and frequently matches or surpasses fully supervised baselines without requiring task-specific training data.

### v2

@@ Abstract
Deep supervised learning models require high volume of labeled data to attain sufficiently good results. Although, the practice of gathering and annotating such big data is costly and laborious. Recently, the application of self supervised learning (SSL) in vision tasks has gained significant attention. The intuition behind SSL is to exploit the synchronous relationships within the data as a form of self-supervision, which can be versatile. In the current big data era, most of the data is unlabeled, and the success of SSL thus relies in finding ways to utilize this vast amount of unlabeled data available. Thus it is better for deep learning algorithms to reduce reliance on human supervision and instead focus on self-supervision based on the inherent relationships within the data. With the advent of ViTs, which have achieved remarkable results in computer vision, it is crucial to explore and understand the various SSL mechanisms employed for training these models specifically in scenarios where there is limited labelled data available. In this survey, we develop a comprehensive taxonomy of systematically classifying the SSL techniques based upon their representations and pre-training tasks being applied. Additionally, we discuss the motivations behind SSL, review popular pre-training tasks, and highlight the challenges and advancements in this field. Furthermore, we present a comparative analysis of different SSL methods, evaluate their strengths and limitations, and identify potential avenues for future research.
@@ 1 Introduction
To tackle above-mentioned challenges in supervised learning, several learning strategies are presented, such as active learning [6], semi-supervised learning [7], and SSL [8]. Recently, Transformers [9] have emerged as an effective Neural Network (NN) architecture for leveraging deeper insights in the input data, specifically, the ViTs [10] are considered to be the alternative of CNNs [11]. ViTs have been successfully employed in tasks related to computer vision, including object detection [12], recognition [13] and semantic segmentation [14]. Notably, ViTs have demonstrated superior performance in image classification, when trained with large volume datasets such as JFT-300M [15].
In essence, by designing pretext tasks, a pre-training phase of network that encourage the model to learn significant patterns in the data itself, SSL methods enable ViTs to understand underlying structures of data without any supervision. The motivation for applying SSL in ViTs is two-fold. Firstly, SSL enables ViTs to employ the vast amount of unlabeled data that is readily available, allowing them to develop more robust and generalizable representations. SSL is beneficial in situations where labeled data is scarce or costly to acquire. Secondly, SSL provides a pathway for pre-training ViTs on larger volume of datasets, followed by of fine-tuning on downstream tasks, leading to improved performance and faster convergence [5].
@@ 1.2 Survey Structure
The paper begins with an introduction to SSL and a detailed discussion on importance of SSL in ViTs. The reported SSL methods are segmented into 5 classes according to their unique features. The paper discuss the recent advancements and applications of SSL across different computer vision tasks. It also presents a taxonomy for SSL, which classifies them based on how feature learning is applied within their architecture. According to this taxonomy, SSLs are divided into five groups, each representing unique way of making use of input features. Frequently used abbreviations are listed in Table 1. The structure of paper is depicted in Fig 1. Section 1 discuss a systematic understanding of the SSL architecture, highlighting its need in ViTs and outlined the advent of SSL architectures. Proceeding to section 2 which covers the advancements in SSL variants, while section 3 provide a taxonomy of the recent SSL architectures, respectively.
Section 4 focuses on the regularization techniques used in SSL, particularly in the area of computer vision. Section 5 presents various metrics and benchmarks for evaluating the effectiveness of SSL, while Section 6 provides a comparative analysis of state-of-the-art SSL mechanisms employing Vision Transformers. Section 7 highlights some open challenges in the domain of SSL, and Section 8 discusses different applications and future directions for SSL in Vision Transformers, including current challenges and potential developments. Finally, Section 9 concludes this survey.
@@
CLIP Contrastive Language-Image pre-training
@@ 2 Evolution of SSL
In computer vision SSL methods are generally segmented into contrastive, generative, and predictive approaches. Contrastive methods, such as MoCo [21] and SimCLR [22], aim to learn patterns by contrasting positive and negative samples. Generative methods, such as GANs [23] and VAEs [24], focus on learning representations by generating samples from the learned representation space. Predictive methods, such as BERT [25] and T5 [26], learn representations by predicting some parts of the input data.
One SSL category is instance discrimination, in which the motto of learning is basically focused on learning patterns by differentiating each sample image from others. This method is rather challenging for large volume of datasets. Examples are SimCLR, SwAV, BYOL, MoCO [22, 27, 28, 21].
@@ 2.1 Evolution of ViTs
In computer vision, the traditional approach in image processing tasks is the utilization of convolutional neural networks (CNNs). However, the CNNs have limitations in capturing dependencies across long spatial distances and global context, both of which are crucial for some vision tasks. ViTs address these limitations via self-attention mechanisms to learn and understand the global context and also dependencies over long spatial distances in the input data .
The initial vision transformer architecture ViT was presented in the paper ”An Image Is Worth 16x16 Words” [10] Since then, various variants of ViTs are proposed, including DeiT [29], Swin Transformer [30], and TNT [31]. These architectures have obtained state-of-the-art performance in different computer vision based tasks, including image classification, object segmentation and object detection.
@@ 2.2 Importance of SSL in ViTs
SSL is especially crucial in the context of ViTs, as it empowers models to learn representations from large amount of unlabeled data. ViTs need vast amounts of data to learn effective representations, and SSL provides a pathway for pre-training on large-scale datasets. By pre-training on large-scale datasets, ViTs can learn robust representations.
@@ 2.3 Evolution of SSL in Computer Vision
SSL technique is used for representation learning where a model is designed to learn the underlying relationship between its inputs. The framework of Energy-Based Models (EBMs) [32] can effectively describe this objective, where the aim of assigning high energy to inputs that are incompatible and low energy to those that are compatible. The foundation for many self-supervised learning techniques applied to images was laid by Geoffrey Hinton and others in the 1980s through the development of autoencoders and the proposal of greedy layer-wise pre-training [33], where layers of a deep neural networks are trained sequentially. However, autoencoders were basically used for learning features and reducing dimensionality. They were impactful at that time as they made it possible to train the first “deep” networks. An analogous approach was Restricted Boltzmann Machines (RBMs), that allows layer-wise training and then combine those layers to build deep belief nets [34]. Although these techniques were eventually replaced by simpler initialization strategies and extended training procedures, they played a vital part in the development of early deep learning architectures.
@@
In 2000s, types of unsupervised learning, including sparse coding and K-means, were employed for self-supervised learning (SSL). These techniques aimed to learn image features without labeled data, laying the groundwork for more complex self-supervised methods. The category of Spatial Context prediction gained attention in SSL for the training of CNNs. In this regard, Pathak et al. (2016) [35] proposed inpainting using context prediction as a pretext task for SSL in images, where the model predicts missing parts of an image. Zhang et al. (2016) [36] explored a straightforward method of the colorization task for self-supervised representation learning. They trained a model to colorize grayscale images, which significantly enhances its capability to understand textures, patterns, and object identities. Noroozi and Favaro (2016) [37] introduced a method where the model learns to solve jigsaw puzzles, using the relative positions of image patches as supervision. Gidaris et al. (2018) [38] introduced RotaNet, which employs a pretext task where the model predicts the rotation angle applied to an image, enhancing feature learning for various orientations. Meanwhile, clustering approaches also progressed in SSL. Caron et al. (2018) [39] proposed DeepCluster [39], a method that iteratively clusters image features and trains a neural network to predict the cluster assignment.
@@
Contrastive learning became significant approach in self-supervised learning during the era of convolutional neural networks (CNNs) [21, 22]. The roots of contrastive learning can be found in the early work on metric learning and the development of Siamese networks [40] in the 1990s and early 2000s. Siamese Network architecture, introduced by Bromley et al. in 1993 [40], involved training twin networks to lessen the relationship among embeddings of matching pairs and expanding the gap between embeddings of constrasting pairs. The formalization of contrastive loss by Hadsell, Chopra, and LeCun (2006) [41] was a significant step forward.
@@
In the late 2010s, the concept of employing contrastive learning for self-supervised learning gained attraction. This period saw the development of several influential techniques that leveraged contrastive objectives to learn meaningful patterns from unlabeled data. Dosovitskiy (2016) [42] used instance discrimination as a pretext task. Each image was handled as a different class, and the network was trained to differentiate between differnt variations of the same image. Wu (2018). [43] refined the instance discrimination approach using a memory repository to save negative examples for the contrastive loss. Oord (2018). [44] introduced method known as Contrastive Predictive Coding in 2018, which learns patterns by predicting future observations in latent space using contrastive learning. Chen (2020) [22] introduced a framework known as SimCLR for contrastive learning of visual features that trains by contrasting similar and non-similar pairs of data. This significantly advanced the performance of self-supervised methods in computer vision. Over time, the idea of hybrid learning gained strength, and in 2020, Caron (2020) [27] introduced SwAV, which combines contrastive learning with clustering by using swapped assignments between different augmentations of an image.
@@
The idea of self-distillation give a new dimension to SSL. Grill (2020) [28] boosted the idea of SSL by introducing BYOL, a self-supervised method that avoids the need for negative samples (simplifying contrastive learning) by using two networks to predict each other’s outputs. He et al. [21] introduced MoCo, which extends the idea of SimCLR by using a dynamic dictionary and a momentum encoder for contrastive learning. Facebook AI Research (Caron [45]) presented DINO, which uses self-distillation with no labels, leveraging VITs for SSL.
@@
In 2020, the idea of transformers in vision was introduced. Vision transformers are high-capacity models and need a huge data for model tuning and good generalization. For improving the learning ability of the transformers, Dosovitskiy (2020) [10] adapted SSL techniques like masked image modeling and contrastive learning for ViT. Inspired by BERT (2018) [25], Dosovitskiy et al. [10] masked certain segment tokens and replaced them with trained mask tokens. Later the model was further trained for predicting pixel values directly. However, they found unsupervised pre-training technique notably less useful than pre-training with supervised methods. Before ViTs, the same idea of inpainting was employed in CNNs, where portions of an image were masked out and the model was taught to reconstruct them, this approach to pre-training is termed as MIM (Masked Image Modeling). In ViTs, the MIM was framed as a regression task. This involved initially utilizing an autoencoder, which encode image segments to separate tokens, subsequently, the transformer is pre-trained to estimate the distinct token values for masked tokens. At that time, Bert pre-training of image transformers (BEiT) [46] demonstrated substantial enhanced results in image classification and semantic segmentation compared to earlier supervised and self-supervised methods. However, its training process was intricate as it required a robust autoencoder to transform image patches into discrete tokens.
@@
Recently, the idea of multi-modality gained attention with the introduction of CLIP by OpenAI in 2021 [51]. CLIP learns image and text embeddings simultaneously by matching corresponding image-caption pairs. Similarly, ALIGN [52] uses self-supervised learning approaches to develop a shared feature space for images and their corresponding captions. The concept of MIM pre-training was streamlined by two concurrent works: MAE by He et al. [53] and SimMIM by Xie et al. [54]. These approaches propose simplified algorithms that immediately rebuild masked parts of image patches instead of using separate image pieces obtained from an encoder as done in BEiT [46]. This idea was inspired by masked language modeling in NLP. These straightforward pre-training strategies performs better as compared to BEiT on classifying images, object detection and semantic segmentation (downstream tasks). MIM has recently dominated this domain by obtaining state-of-the-art performance on ViTs. The essence of this technique is to enhance the network’s capability to capture visual context at patch level using a denoising auto-encoding mechanism. Invariant Joint Embedding Predictive Architecture (IJEPA) [55] has emerged as a promising alternative, addressing some of the limitations inherent in other self-supervised learning methods. Unlike traditional self-supervised approaches that depend heavily on hand-crafted data augmentations, IJEPA focuses on invariant feature learning and joint embeddings. IJEPA relies on predicting the patterns of destination blocks within an image from a single context block, using a strategic masking method to guide the learning process.
@@
RotNet [38] Spatial awareness, Object recognition Complex texture understanding
@@ 2.4 Comparison of SSL with Transfer Learning
In computer vision, transfer learning (TL) [56] is employed in the development of CNN models to solve a new challenging problem by utilizing the pre-trained model’s weight sharing mechanism especially in the presence of data scarcity [57]. Initially, transfer learning was the dominant approach, but recently, SSL has demonstrated promising results across various applications.
@@ 3.0.1 Contrastive Methods
Contrastive Learning (CL) stands as a prominent technique in SSL [21, 22]. Its aim is to capture invariant semantics through pairs of random views, termed Contrastive multi-view coding. CL ensures that global projections of representations are similar for positive samples and dissimilar for negative samples. This section delves into literature reviews on Contrastive Learning applied in transformer frameworks.
@@
Wang and co-workers in 2022 [64] introduced the method SRCL that is semantically-relevant contrastive learning for histopathology image analysis. Unlike traditional contrastive learning, SRCL evaluates similarity among instances to identify additional positive pairs by arranging various positive examples that share related visual concepts. This approach enriches the diversity of positive pairs and yields more informative patterns. They employee a hybrid model named CTransPath, in which they combines a CNN with a multi-scale Swin Transformer architecture, their methodology acts as a unified feature extractor, addressing both local and global features during pre-training, seeking to acquire global feature interpretations tailored for activities within the domain of histopathology images. The efficacy of SRCL-pretrained CTransPath was evaluated across five distinct downstream tasks spanning nine publicly available datasets, such as retrieving patches, classifying patches, performing weakly-supervised classification of whole-slide images, detecting mitosis, and segmenting glands in colorectal adenocarcinoma. The findings demonstrated that the visual representations derived from SRCL not just achieved state-of-the-art performance throughout all datasets but it also exhibited increased robustness and transferability compared to both supervised and self-supervised ImageNet pre-training techniques.
@@
To address instability issues during training, Chen and his colleagues in 2021 [65] utilized SSL framework in MOCO, representing an advancement in Contrastive Learning (CL) methodologies. MOCO utilizes dictionaries to gauge similarity between encoded keys and queries. Empirical assessments suggested that self-supervised Transformers within the MOCO framework can demonstrate good performance with relatively fewer inductive biases compared to ImageNet supervised ViT. This approach aims for a better balance between simplicity, accuracy, and scalability, showcasing its superiority over supervised pre-training methods in multiple detection and segmentation tasks. The study proposes replacing the patch projection layer in ViT with fixed random patch projections and suggests removing positional embeddings in SSL setups to address instability.
@@
Mo and his team [66] in 2023 introduced MCVT, which focuses on projecting class tokens to embedding space during the initial or final phases of the ViTs using multi-layer perceptrons. In this methodology, InfoNCE loss is applied to low-level features while ProtoNCE loss is introduced in high-level features, enhancing performance in downstream classification tasks. Extensive experimentation transferring MCVT pre-trained backbones to various downstream tasks confirmed its efficacy across multiple vision-related tasks.
@@
Radford in 2021 [51] proposed a Contrastive Language-Image pre-training (CLIP) framework, which incorporates transformers, learns image representations through natural language supervision. In pre-training, CLIP concurrently trains image and text encoders to accurately predict corresponding pairs within a batch of (image, text) training instances. CLIP discerns among the N * N possible (image, text) pairings within a batch. It achieves this by learning a multi-modal embedding space, optimizing the image and text encoders jointly to increase the cosine similarity of embeddings for the N genuine pairs while decreasing the resemblance of embeddings for the ( N 2̂ - N ) incorrect pairings. Symmetric cross-entropy loss is used in this optimization strategy to adjust these similarity scores. After pre-training, the model uses natural language to reference acquired visual concepts, facilitating zero-shot transfer to downstream tasks, demonstrating substantial transferability to diverse computer vision tasks without the need for task-specific training data. Evaluation encompasses benchmarking across more than 30 diverse computer vision datasets. Remarkably, the model demonstrates substantial transferability to most tasks and frequently matches or surpasses fully supervised baselines without requiring task-specific training data.